# Uniform approximation on certain polynomial polyhedra in $\mathbb{C}^{2}$

### v1

@@ Abstract.
In this paper we extend the dichotomy given by Samuelsson and Wold [28] that can be thought of as an analogue of the Wermer maximality theorem in $\mathbb{C}^{2}$ for certain polynomial polyhedra. We consider complex non-degenerate simply connected polynomial polyhedra of the form $\Omega:=\{z\in\mathbb{C}^{2}:|p_{1}(z)|<1,|p_{2}(z)|<1\}$ such that $\overline{\Omega}$ is compact. Under a mild condition of the polynomials $p_{1}$ and $p_{2}$, we prove that either the uniform algebra, generated by polynomials and some continuous functions $f_{1},\dots,f_{N}$ on the distinguished boundary that extends as pluriharmonic functions on $\Omega$, is all continuous functions on the distinguished boundary or there exists an algebraic variety in $\overline{\Omega}$ on which each $f_{j}$ is holomorphic. We also compute the polynomial hull of the graph of pluriharmonic functions in some cases where the pluriharmonic functions are conjugates of holomorphic polynomials. We also give a couple of general theorem about uniform approximation on the domains with low boundary regularity.
@@ 1. Introduction and statements of the results
Result 1.1 uses the properties of the bidisc crucially. One can ask the question: Does the same dichotomy hold when one considers other domains? This does not have immediate answer. For some images of the bidisc under proper polynomial maps, symmetrized bidisc being one of them, the question is answered positively in [11]. In this paper, we prove a similar result for certain polynomial polyhedra. Let $p_{1},p_{2},\cdots,p_{l}$ be holomorphic polynomials in $z_{1},z_{2},\ldots,z_{n}$. Consider the set $\mathfrak{D}_{l}$ known as the polynomial polyhedron defined by
@@
is a proper map. Let $P$ be a holomorphic polynomial. We will compute the polynomially convex hull of the graph of $\overline{P\circ\Psi}$ over the distinguished boundary of the polynomial polyhedron $\mathfrak{D}_{2}$ (see Section 4 for details). Jimbo [18, 19] gave the description of polynomial hulls of graphs of anti-holomorphic polynomials on the torus. Here, we treat the case when $f$ is the restriction of an anti-holomorphic polynomial in $\mathbb{C}^{2}$ to the distinguished boundary $\Gamma_{{\mathfrak{D}_{2}}}$ of a complex non-degenerate polynomial polyhedron $\mathfrak{D}_{2}$ in $\mathbb{C}^{2}.$ We present explicit descriptions of the hull of the graph of $f$ on $\Gamma_{{\mathfrak{D}_{2}}}.$ In this setting, we show that the set $\widehat{{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})}\setminus{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})$ is contained in an algebraic variety. We now look at the graph over certain bounded domains whose closure is polynomially convex. Specifically, we look at domains with polynomially convex closure and having segment property.
@@
In [4] authors showed that a domain $\Omega$ has the segment property if, and only if, $\Omega$ has $\mathcal{C}^{0}$ boundary. For more details on the segment property we refer the reader to [7]. We now present a couple of observations that will generalize results from ([28, Theorem 1.3]) and [17, Theorem 1.1] respectively:
@@
We observe that the $\mathcal{C}^{1}$-smoothness of $\Omega$ is required to prove ([28, Theorem 1.3]) and ([17, Theorem 1.1]) only to show that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex. By Lemma 2.21, we can say that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex for any bounded domain $\Omega$ with polynomially convex closure and having segment property except possibly at a countable set of boundary points. The rest follows from [28, Theorem 1.3] and [17, Theorem 1.1], respectively. We now give a brief outline for our proofs of Theorem 1.2. The philosophy behind the proof of Theorem 1.2 closely follows that of Result 1.1 although adapting the relevant tools to our case requires some new insights. For example, in the proof of Theorem 1.2, the polynomial convexity of $h$ over the polynomial polyhedron is crucial, which we prove by showing that the polynomial polyhedron has segment property. We also gave a version of Tornehave’s result for polynomial polyhedra. Firstly, we establish that Theorem 1.2 holds if and only if the graph of $h$ over the distinguished boundary of the polynomial polyhedron $\mathfrak{D}_{2}$ is polynomially convex (refer to Theorem 2.24). Therefore, if the graph lacks polynomial convexity, we examine different scenarios through specific lemmas, demonstrating the existence of an analytic variety within $\mathfrak{D}_{2}$ where all $h_{j}$’s are holomorphic. Utilizing Tornehave’s result (for polynomial polyhedron), we can find the required algebraic variety. A few words about the layout of the paper: The first part of Section 2 collects some known results that are required in this paper while in the second part we prove several new results which are crucial for this paper. In Section 3 we present a proof of Theorem 1.2. The polynomial hulls of some graphs are computed in Section 4. Section 5 is devoted to some examples.
@@ 2. Technical Results
We make use of the following result due to Samuelsson-Wold [28, Corollary 4.2].
@@
([28, Propostion 4.7]) Let $K\subset\mathbb{C}^{N}$ be compact, let $F:\mathbb{C}^{n}\to\mathbb{C}^{m}$ be the uniform limit on $K$ of entire functions, and let $Y=F(K);$ note that $F$ extends to $\widehat{K}.$ If $y\in Y$ is a peak point for the algebra $\mathscr{P}(Y),$ then $\widehat{K}\cap F^{-1}(y)=\widehat{(K\cap F^{-1}\{y\}}).$
Let ${\sf psh}(\Omega)$ denote the class of plurisubharmonic functions in $\Omega$ and ${\sf psh}(\overline{\Omega})$ denotes the collection of function $\phi$ such that there exists an open set $U\ \supset\Omega$ with $\phi\in{\sf psh}(U)$. Sibony [29, Theorem 2.2] showed that if $\Omega$ is a smooth pseudoconvex domain, then every functions in ${\sf psh}(\Omega)\cap C(\overline{\Omega})$ can be approximated uniformly with functions in ${\sf psh}(\overline{\Omega})\cap C^{\infty}(\overline{\Omega})$ and we say $\Omega$ has the PSH-Mergelyan property. Fornæss and Wiegerinck [9, Theorem 1] proved that this is also true for any bounded $\mathcal{C}^{1}$-smooth domain. Avelin, Hed, and Persson [5] further relaxed the boundary conditions to continuity. Subsequently, in [21], the authors presented an improved version of this result.
@@
The following result [27, Theorem 5.1] gives a criterion for a homogeneous polynomial map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n}$ to be proper holomorphic.
@@
[…]
@@
The following is known as Tornehave’s result ([30, Corollary 3.8.11]).
@@
We denote $\Sigma_{j}:=\{z\in\mathbb{C}^{n}:\psi_{j}(z)=0\}$ and $\Sigma_{J_{k}}:=\Sigma_{j_{1}}\cap\cdots\cap\Sigma_{j_{k}}$, where $J_{k}=({j_{1},\cdots,j_{k}})$ with $1\leq j_{1}<j_{2}<\ldots<j_{k}\leq N$. The topological boundary of $\Omega$ is denoted by $\partial\Omega$ and clearly $\partial\Omega\subset\cup_{j}\Sigma_{j}.$ The result presented below describes a class of domains where the segment property holds. A special case for Weil polyhedra was proved in [22, Proposition 2.1].
@@
The proof of the following result for a $C^{1}$-smooth domain is given in Samuelsson and Wold [28, Lemma 4.5].
@@
Since $S\subset V_{Sing}$ and the set of regular points of an irreducible variety is connected, therefore we can say that $V\setminus S$ is connected. But right hand side of (2.13) is disconnected, and this is a contradiction to the assumption that $V\nsubseteq Z_{1},V\nsubseteq Z_{2}$ but $V\subset Z_{1}\cup Z_{2}.$ Without loss of generality, we assume that $V\subset Z_{1}.$ Next, we claim that $V=\mathfrak{D}_{n}\cap Z_{1}.$ Assume that $V\subset Z_{1}\cap\mathfrak{D}_{n}.$ Since $Z_{1}$ is an irreducible variety and if $V$ is proper subset of $Z_{1},$ then $\dim_{\mathbb{C}}V<\dim_{\mathbb{C}}Z_{1}$ (see Izzo [3], Federer [8, Theorem 3.4.8, assertion (15)]). This is not possible, as otherwise, $V$ will be discrete. Therefore, $V=\mathfrak{D}_{n}\cap Z_{1}.$
@@
We claim that $K$ is polynomially convex. To prove this claim, we assume that $(a,b)\in\mathbb{C}^{2n}\setminus K,$ where $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n}).$ Case I: $b_{l}p_{l}(a)\not=1,$ $|p_{l}(a)|=1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=w_{l}p_{l}(z)-1.$ Then
@@
Hence, in this case, $(a,b)\notin\widehat{K}.$ Case II: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|<1$ for some $l\in\{1,\cdots,n\}.$ This implies $|b_{l}|>1.$ Consider the polynomial $\psi(z,w)=w_{l}.$ Then
@@
Hence, in this case, $(a,b)\notin\widehat{K}.$ Case III: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|>1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=p_{l}(z).$ Then
@@
Hence, in this case, $(a,b)\notin\widehat{K}.$ Combining all three cases, we conclude that $K$ is polynomially convex. Since $dp_{1}\wedge\cdots\wedge dp_{n}(z)\neq 0\text{ on }\Gamma,$ $\Gamma$ is totally real, and being graph over a totally real manifold, $K$ is also totally real. Therefore, we get that $\mathscr{P}(K)=\mathcal{C}(K).$ Since $K$ is the graph of $(\overline{p_{1}},\cdots,\overline{p_{n}})$ on $\Gamma,$ hence
@@
Now we will prove the main result. Assume $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma).$ Define $X={\sf Gr}_{f}(\Gamma).$ Since $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma),$ then $\mathscr{P}(X)=\mathcal{C}(X).$ Hence $X$ is polynomially convex. Conversely, assume that ${\sf Gr}_{f}(\Gamma)$ is polynomially convex. Let $\Omega$ be a neighborhood of $\Gamma$ such that $p_{1}(z)\not=0~{}~{}\text{ on }\Omega.$ Define $\phi(z,w)=\frac{1}{p_{1}(z)}.$ Then $\phi$ is holomorphic on $\Omega.$ Also, $\phi$ is holomorphic on $\Omega\times\mathbb{C}^{n}.$ Since ${\sf Gr}_{f}(\Gamma)\subset\Omega\times\mathbb{C}^{n},$ by the Oka-Weil approximation theorem, there exists a sequence of polynomials $P_{j}$ in $\mathbb{C}^{n}_{z}\times\mathbb{C}^{n}_{w}$ such that $P_{j}\to\phi$ uniformly on ${\sf Gr}_{f}(\Gamma).$ This implies $P_{j}(z,f(z))\to\phi(z,w)=\frac{1}{p_{1}(z)}=\overline{p_{1}(z)}$ uniformly on $\Gamma.$ Hence $\overline{p_{1}(z)}\in[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$ By the similar method we can show that $\overline{{p_{j}}}\in[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]~{}\forall_{j}\in\{1,\cdots,n\}.$ Hence, $[z_{1},\cdots,z_{n},\overline{p_{1}},\cdots,\overline{p_{n}};{\Gamma}]\subset[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$ Using (2.14), we get that
@@
Let $h_{j}$ denote the pluriharmonic extension of $f_{j}$ to $\mathfrak{D}_{2},$ and write $h=(h_{1},\cdots,h_{N}):\overline{\mathfrak{D}}_{2}\to\mathbb{C}^{N}.$ The following two cases may occur: Case I: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is polynomially convex. In view of Theorem 2.24, we have that
@@
Therefore, in this case, the proof of the theorem is complete. Case II: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex. By Lemma 2.22, it is enough to find an irreducible analytic variety $V\subset\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$ Since ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, there exists $\alpha\in\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Now, we discuss different situations through some lemmas.
@@
Let $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}).$ We will find $\max|\omega|$ on $\Gamma_{\mathfrak{D}_{2}},$ i.e., we need to find $\max|\omega|$ on $|\omega-p_{1}(\alpha)|=|p_{1}(z)|=1.$ Let $\omega=x+iy~{}~{}\text{ and }p_{1}(\alpha)=\lambda_{1}+i\lambda_{2}.$ The equation of the straight line through the origin and through the centre of the circle $|\omega-p_{1}(\alpha)|=1$ is given by $y=\frac{\lambda_{2}}{\lambda_{1}}x.$ Solving $y=\frac{\lambda_{2}}{\lambda_{1}}x$ and $(x-\lambda_{1})^{2}+(y-\lambda_{2})^{2}=1,$ we get that $(x,y)=(0,0),~{}~{}\text{ or }(x,y)=(2\lambda_{1},2\lambda_{2}).$ Therefore, $\max|\omega|=2$ and maximum attains at $\omega=2\lambda_{1}+i2\lambda_{2}=2p_{1}(\alpha).$ Then $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}),$ implies $p_{1}(z)=p_{1}(\alpha).$ Therefore, $\max|\psi|=1$ on $\Gamma_{\mathfrak{D}_{2}}$ and since $|\omega|<1$ if $\omega\not=2p_{1}(\alpha),$ $|\psi|<1\text{ on }\Gamma_{\mathfrak{D}_{2}}\setminus\partial\triangle_{\alpha}.$ Finally we define the polynomial in $\mathbb{C}^{2+N}$
@@
Because of Result 2.10 and using the fact that the finite intersection of analytic varieties is again an analytic variety, we get that $\widetilde{V}$ is an analytic variety in $\mathfrak{D}_{2}$. By Result 2.1, we get that ${\sf Gr}_{h}(\mathfrak{D}_{2}\setminus\widetilde{V})$ is totally real. Let $E={\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ and $\Omega=(\mathfrak{D}_{2}\setminus\widetilde{V})\times\mathbb{C}^{N}.$ Then

### v2

@@ Abstract.
In this paper we extend the dichotomy given by Samuelsson and Wold that can be thought of as an analogue of the Wermer maximality theorem in $\mathbb{C}^{2}$ for certain polynomial polyhedra. We consider complex non-degenerate simply connected polynomial polyhedra of the form $\Omega:=\{z\in\mathbb{C}^{2}:|p_{1}(z)|<1,|p_{2}(z)|<1\}$ such that $\overline{\Omega}$ is compact. Under a mild condition of the polynomials $p_{1}$ and $p_{2}$, we prove that either the uniform algebra, generated by polynomials and some continuous functions $f_{1},\dots,f_{N}$ on the distinguished boundary that extends as pluriharmonic functions on $\Omega$, is all continuous functions on the distinguished boundary or there exists an algebraic variety in $\overline{\Omega}$ on which each $f_{j}$ is holomorphic. We also compute the polynomial hull of the graph of pluriharmonic functions in some cases where the pluriharmonic functions are conjugates of holomorphic polynomials. We also give a couple of general theorem about uniform approximation on the domains with low boundary regularity.
@@ 1. Introduction and statements of the results
Result 1.1 uses the properties of the bidisc crucially. One can ask the question: Does the same dichotomy hold when one considers other domains? This does not have immediate answer. For some images of the bidisc under proper polynomial maps, symmetrized bidisc being one of them, the question is answered positively in [11].
In this paper, we prove a similar result for certain polynomial polyhedra. Let $p_{1},p_{2},\cdots,p_{l}$ be holomorphic polynomials in $z_{1},z_{2},\ldots,z_{n}$. Consider the set $\mathfrak{D}_{l}$ known as the polynomial polyhedron defined by
@@
is a proper map. Let $P$ be a holomorphic polynomial. We will compute the polynomially convex hull of the graph of $\overline{P\circ\Psi}$ over the distinguished boundary of the polynomial polyhedron $\mathfrak{D}_{2}$ (see Section 4 for details). Jimbo [18, 19] gave the description of polynomial hulls of graphs of anti-holomorphic polynomials on the torus. Here, we treat the case when $f$ is the restriction of an anti-holomorphic polynomial in $\mathbb{C}^{2}$ to the distinguished boundary $\Gamma_{{\mathfrak{D}_{2}}}$ of a complex non-degenerate polynomial polyhedron $\mathfrak{D}_{2}$ in $\mathbb{C}^{2}.$ We present explicit descriptions of the hull of the graph of $f$ on $\Gamma_{{\mathfrak{D}_{2}}}.$ In this setting, we show that the set $\widehat{{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})}\setminus{\sf Gr}_{f}(\Gamma_{{\mathfrak{D}_{2}}})$ is contained in an algebraic variety.
We now look at the graph over certain bounded domains whose closure is polynomially convex. Specifically, we look at domains with polynomially convex closure and having segment property.
@@
In [4] authors showed that a domain $\Omega$ has the segment property if, and only if, $\Omega$ has $\mathcal{C}^{0}$ boundary. For more details on the segment property we refer the reader to [7]. We now present a couple of observations that will generalize results from ([30, Theorem 1.3]) and [17, Theorem 1.1] respectively:
@@
We observe that, in ([17, Theorem 1.1]) and ([30, Theorem 1.3]), the $\mathcal{C}^{1}$-smoothness of $\Omega$ is required to prove and only to show that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex. By Lemma 2.21, we can say that ${\sf Gr}_{h}(\overline{\Omega})$ is polynomially convex for any bounded domain $\Omega$ with polynomially convex closure and having segment property except possibly at a countable set of boundary points. The rest follows from [30, Theorem 1.3] and [17, Theorem 1.1], respectively.
Although the proof of Theorem 1.6 and Theorem 1.7 are not very different from that of [17] except for the polynomial convexity, but there are huge implication of Theorem 1.6 and Theorem 1.7 as we can deal with a huge class of nonsmooth domains here. A domain having continuous boundary satisfies the segment property. Although whether the closure of a bounded domain is polynomially convex is a very difficult problem, the class of bounded convex domains satisfies the segment property and their closure is also polynomially convex. Hence, we obtain the following corollary:
Corollary 1.8.
Let $\Omega$ be a bounded convex domain in $\mathbb{C}^{n}$. Let $h_{j}$, $j=1,\dots,N$, be pluriharmonic in $\Omega$ and continuous on $\overline{\Omega}$. Assume further that there does not exist a nontrivial analytic disc $\triangle\subset\Omega$ on which all the $h_{j}$’s are holomorphic. Then
  1. (1)
    $G_{h}(\partial\Omega)$ is polynomially convex.
  2. (2)
  3. (3)
We now provide some remarks: the first two of them concerns about the comparison between Theorem 1.2 with known results on polynomial polyhedra and other domains in $\mathbb{C}^{2}$; and the other concerns about the possibility of extending Theorem 1.2 to domains in $\mathbb{C}^{n}$, $n\geq 3$. These comments might be interesting in the context of further research.
Remark 1.9.
The conditions on polynomial polyhedra in Theorem 1.2 are quite natural. On simply connected domain any complex valued pluriharmonic function can be seen as a sum of a holomorphic function and the real part of another holomorphic function. This is used crucially in our proofs. Complex non-degeneracy of the polyhedra gives us the segment property which is one other essential ingredient in our proof. A natural class of polynomial polyhedra is special polynomial polyhedra in the sense of [21]. The class of polyhedra that satisfies conditions of Theorem 1.2 has nontrivial intersection with the class of special polynomial polyhedra. Any polynomially convex set satisfying some conditions (see Corollary 2 and Theorem 3 of [21]) can be approximated by special polynomial polyhedra. It is not clear whether Theorem 1.2 holds for special polynomial polyhedra defined in [21]. Also, there are some approximation results of Mergelyan type on certain polynomial polyhedra (see [23], [24]).
Remark 1.10.
When one considers a general domain in $\mathbb{C}^{2}$, it is already difficult to find the Shilov boundary. Therefore, one assumes the boundary approximation as in Theorem 1.6. If Shilov boundary can be found precisely and it is two real dimension and the other parts of the boundary are foliated by analytic discs, then the question of getting a dichotomy as in Theorem 1.2 make sense and is also interesting. For certain proper holomorphic images of the bidisc such dichotomy were shown in [11].
Remark 1.11.
In complex dimension three or higher, even for the polydisc, things are complicated. Main obstruction comes in giving a stratification of the graph over the boundary. This is the reason for assuming boundary approximation while considering the general domains in Theorem 1.6. Here, unlike the case in Theorem 1.2, one need to deal with $n-1$ dimensional analytic varieties; some part of our techniques fail even in $\mathbb{C}^{3}$. The case of polynomial polyhedra could be thought of as the first step towards it.
@@
We now give a brief outline for our proofs of Theorem 1.2. The philosophy behind the proof of Theorem 1.2 closely follows that of Result 1.1 although adapting the relevant tools to our case requires some new insights. For example, in the proof of Theorem 1.2, the polynomial convexity of $h$ over the polynomial polyhedron is crucial, which we prove by showing that the polynomial polyhedron has segment property. We also gave a version of Tornehave’s result for polynomial polyhedra. Firstly, we establish that Theorem 1.2 holds if and only if the graph of $h$ over the distinguished boundary of the polynomial polyhedron $\mathfrak{D}_{2}$ is polynomially convex (refer to Theorem 2.24). Therefore, if the graph lacks polynomial convexity, we examine different scenarios through specific lemmas, demonstrating the existence of an analytic variety within $\mathfrak{D}_{2}$ where all $h_{j}$’s are holomorphic. Utilizing Tornehave’s result (for polynomial polyhedron), we can find the required algebraic variety.
A few words about the layout of the paper: The first part of Section 2 collects some known results that are required in this paper while in the second part we prove several new results which are crucial for this paper. In Section 3 we present a proof of Theorem 1.2. The polynomial hulls of some graphs are computed in Section 4. Section 5 is devoted to some examples.
@@ 2. Technical Results
We make use of the following result due to Samuelsson-Wold [30, Corollary 4.2].
@@
([30, Propostion 4.7]) Let $K\subset\mathbb{C}^{N}$ be compact, let $F:\mathbb{C}^{n}\to\mathbb{C}^{m}$ be the uniform limit on $K$ of entire functions, and let $Y=F(K);$ note that $F$ extends to $\widehat{K}.$ If $y\in Y$ is a peak point for the algebra $\mathscr{P}(Y),$ then $\widehat{K}\cap F^{-1}(y)=\widehat{(K\cap F^{-1}\{y\}}).$
Let ${\sf psh}(\Omega)$ denote the class of plurisubharmonic functions in $\Omega$ and ${\sf psh}(\overline{\Omega})$ denotes the collection of function $\phi$ such that there exists an open set $U\ \supset\Omega$ with $\phi\in{\sf psh}(U)$. Sibony [31, Theorem 2.2] showed that if $\Omega$ is a smooth pseudoconvex domain, then every functions in ${\sf psh}(\Omega)\cap C(\overline{\Omega})$ can be approximated uniformly with functions in ${\sf psh}(\overline{\Omega})\cap C^{\infty}(\overline{\Omega})$ and we say $\Omega$ has the PSH-Mergelyan property. Fornæss and Wiegerinck [9, Theorem 1] proved that this is also true for any bounded $\mathcal{C}^{1}$-smooth domain. Avelin, Hed, and Persson [5] further relaxed the boundary conditions to continuity. Subsequently, in [22], the authors presented an improved version of this result.
@@
The following result [29, Theorem 5.1] gives a criterion for a homogeneous polynomial map from $\mathbb{C}^{n}$ to $\mathbb{C}^{n}$ to be proper holomorphic.
@@
$Z^{\prime}:=\{z\in Z_{\text{reg}}:i^{*}_{Z_{\text{reg}}}(\overline{\partial}h_{j_{1}}\wedge\cdots\wedge\overline{\partial}h_{j_{n}})(z)=0\}.$
@@
The following is known as Tornehave’s result ([32, Corollary 3.8.11]).
@@
We denote $\Sigma_{j}:=\{z\in\mathbb{C}^{n}:\psi_{j}(z)=0\}$ and $\Sigma_{J_{k}}:=\Sigma_{j_{1}}\cap\cdots\cap\Sigma_{j_{k}}$, where $J_{k}=({j_{1},\cdots,j_{k}})$ with $1\leq j_{1}<j_{2}<\ldots<j_{k}\leq N$. The topological boundary of $\Omega$ is denoted by $\partial\Omega$ and clearly $\partial\Omega\subset\cup_{j}\Sigma_{j}.$
The result presented below describes a class of domains where the segment property holds. A special case for Weil polyhedra was proved in [23, Proposition 2.1].
@@
The proof of the following result for a $C^{1}$-smooth domain is given in Samuelsson and Wold [30, Lemma 4.5].
@@
Since $S\subset V_{Sing}$ and the set of regular points of an irreducible variety is connected, therefore we can say that $V\setminus S$ is connected. But right hand side of (2.10) is disconnected, and this is a contradiction to the assumption that $V\nsubseteq Z_{1},V\nsubseteq Z_{2}$ but $V\subset Z_{1}\cup Z_{2}.$ Without loss of generality, we assume that $V\subset Z_{1}.$
Next, we claim that $V=\mathfrak{D}_{n}\cap Z_{1}.$ Assume that $V\subset Z_{1}\cap\mathfrak{D}_{n}.$ Since $Z_{1}$ is an irreducible variety and if $V$ is proper subset of $Z_{1},$ then $\dim_{\mathbb{C}}V<\dim_{\mathbb{C}}Z_{1}$ (see Izzo [3], Federer [8, Theorem 3.4.8, assertion (15)]). This is not possible, as otherwise, $V$ will be discrete. Therefore, $V=\mathfrak{D}_{n}\cap Z_{1}.$
@@
We claim that $K$ is polynomially convex. To prove this claim, we assume that $(a,b)\in\mathbb{C}^{2n}\setminus K,$ where $a=(a_{1},\cdots,a_{n})$ and $b=(b_{1},\cdots,b_{n}).$
Case I: $b_{l}p_{l}(a)\not=1,$ $|p_{l}(a)|=1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=w_{l}p_{l}(z)-1.$ Then
@@
Hence, in this case, $(a,b)\notin\widehat{K}.$
Case II: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|<1$ for some $l\in\{1,\cdots,n\}.$ This implies $|b_{l}|>1.$ Consider the polynomial $\psi(z,w)=w_{l}.$ Then
@@
Hence, in this case, $(a,b)\notin\widehat{K}.$
Case III: $b_{l}p_{l}(a)=1,$ $|p_{l}(a)|>1$ for some $l\in\{1,\cdots,n\}.$ Consider the polynomial $\psi(z,w)=p_{l}(z).$ Then
@@
Hence, in this case, $(a,b)\notin\widehat{K}.$
Combining all three cases, we conclude that $K$ is polynomially convex. Since $dp_{1}\wedge\cdots\wedge dp_{n}(z)\neq 0\text{ on }\Gamma,$ $\Gamma$ is totally real, and being graph over a totally real manifold, $K$ is also totally real. Therefore, we get that $\mathscr{P}(K)=\mathcal{C}(K).$ Since $K$ is the graph of $(\overline{p_{1}},\cdots,\overline{p_{n}})$ on $\Gamma,$ hence
@@
Now we will prove the main result. Assume $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma).$ Define $X={\sf Gr}_{f}(\Gamma).$ Since $[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]=\mathcal{C}(\Gamma),$ then $\mathscr{P}(X)=\mathcal{C}(X).$ Hence $X$ is polynomially convex.
Conversely, assume that ${\sf Gr}_{f}(\Gamma)$ is polynomially convex. Let $\Omega$ be a neighborhood of $\Gamma$ such that $p_{1}(z)\not=0~{}~{}\text{ on }\Omega.$ Define $\phi(z,w)=\frac{1}{p_{1}(z)}.$ Then $\phi$ is holomorphic on $\Omega.$ Also, $\phi$ is holomorphic on $\Omega\times\mathbb{C}^{n}.$ Since ${\sf Gr}_{f}(\Gamma)\subset\Omega\times\mathbb{C}^{n},$ by the Oka-Weil approximation theorem, there exists a sequence of polynomials $P_{j}$ in $\mathbb{C}^{n}_{z}\times\mathbb{C}^{n}_{w}$ such that $P_{j}\to\phi$ uniformly on ${\sf Gr}_{f}(\Gamma).$ This implies $P_{j}(z,f(z))\to\phi(z,w)=\frac{1}{p_{1}(z)}=\overline{p_{1}(z)}$ uniformly on $\Gamma.$ Hence $\overline{p_{1}(z)}\in[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$ By the similar method we can show that $\overline{{p_{j}}}\in[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma]~{}\forall_{j}\in\{1,\cdots,n\}.$ Hence, $[z_{1},\cdots,z_{n},\overline{p_{1}},\cdots,\overline{p_{n}};{\Gamma}]\subset[z_{1},\cdots,z_{n},f_{1},\cdots,f_{N};\Gamma].$ Using (2.11), we get that
@@
Let $h_{j}$ denote the pluriharmonic extension of $f_{j}$ to $\mathfrak{D}_{2},$ and write $h=(h_{1},\cdots,h_{N}):\overline{\mathfrak{D}}_{2}\to\mathbb{C}^{N}.$ The following two cases may occur:
Case I: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is polynomially convex. In view of Theorem 2.24, we have that
@@
Therefore, in this case, the proof of the theorem is complete.
Case II: ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex. By Lemma 2.22, it is enough to find an irreducible analytic variety $V\subset\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that $\overline{V}\setminus V\subset\Gamma_{\mathfrak{D}_{2}}.$
Since ${\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ is not polynomially convex, there exists $\alpha\in\overline{\mathfrak{D}}_{2}\setminus\Gamma_{\mathfrak{D}_{2}}$ such that ${\sf Gr}_{h}(\alpha)\in\widehat{{\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})}.$ Now, we discuss different situations through some lemmas.
@@
Let $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}).$ We will find $\max|\omega|$ on $\Gamma_{\mathfrak{D}_{2}},$ i.e., we need to find $\max|\omega|$ on $|\omega-p_{1}(\alpha)|=|p_{1}(z)|=1.$ Let $\omega=x+iy~{}~{}\text{ and }p_{1}(\alpha)=\lambda_{1}+i\lambda_{2}.$ The equation of the straight line through the origin and through the centre of the circle $|\omega-p_{1}(\alpha)|=1$ is given by $y=\frac{\lambda_{2}}{\lambda_{1}}x.$ Solving $y=\frac{\lambda_{2}}{\lambda_{1}}x$ and $(x-\lambda_{1})^{2}+(y-\lambda_{2})^{2}=1,$ we get that $(x,y)=(0,0),~{}~{}\text{ or }(x,y)=(2\lambda_{1},2\lambda_{2}).$ Therefore, $\max|\omega|=2$ and maximum attains at $\omega=2\lambda_{1}+i2\lambda_{2}=2p_{1}(\alpha).$ Then $\omega=p_{1}(\alpha)+p_{1}(z_{1},z_{2}),$ implies $p_{1}(z)=p_{1}(\alpha).$ Therefore, $\max|\psi|=1$ on $\Gamma_{\mathfrak{D}_{2}}$ and since $|\omega|<1$ if $\omega\not=2p_{1}(\alpha),$ $|\psi|<1\text{ on }\Gamma_{\mathfrak{D}_{2}}\setminus\partial\triangle_{\alpha}.$
Finally we define the polynomial in $\mathbb{C}^{2+N}$
@@
Because of Result 2.10 and using the fact that the finite intersection of analytic varieties is again an analytic variety, we get that $\widetilde{V}$ is an analytic variety in $\mathfrak{D}_{2}$.
By Result 2.1, we get that ${\sf Gr}_{h}(\mathfrak{D}_{2}\setminus\widetilde{V})$ is totally real. Let $E={\sf Gr}_{h}(\Gamma_{\mathfrak{D}_{2}})$ and $\Omega=(\mathfrak{D}_{2}\setminus\widetilde{V})\times\mathbb{C}^{N}.$ Then